OpenStudy (anonymous):

Please help me find the (k+1) statement for: 2 + 4 + 8 + . . . + 2n = 2(n+1) – 2 I already did the 1 and k statements. They are(I think): P1 2(1+1)-2 2(2)-2 4-2=2 P1 = 1 = 2 Pk Pk = 2 + 4 + 8 + … + 2k = 2(k+1) – 2

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

Please Help!

jimthompson5910 (jim_thompson5910):

Pk : 2 + 4 + 8 + … + 2k = 2(k+1) – 2 P(k+1) : 2 + 4 + 8 + … + 2k + 2(k + 1)= 2((k+1)+1) – 2 does that help?

OpenStudy (anonymous):

Thanks!